# Guided Interpretable Facial Expression Recognition via Spatial Action Unit Cues

### v1

@@ Abstract
While state-of-the-art facial expression recognition (FER) classifiers achieve a high level of accuracy, they lack interpretability, an important aspect for end-users. To recognize basic facial expressions, experts resort to a codebook associating a set of spatial action units to a facial expression. In this paper, we follow the same expert footsteps, and propose a learning strategy that allows us to explicitly incorporate spatial action units (AUs) cues into the classifier’s training to build a deep interpretable model. In particular, using this AUs codebook, input image expression label, and facial landmarks, a single action units heatmap is built to indicate the most discriminative regions of interest in the image w.r.t the facial expression. We leverage this valuable spatial cue to train a deep interpretable classifier for FER. This is achieved by constraining the spatial layer features of a classifier to be correlated with AUs map. Using a composite loss, the classifier is trained to correctly classify an image while yielding interpretable visual layer-wise attention correlated with AUs maps, simulating the experts’ decision process. This is achieved using only the image class expression as supervision and without any extra manual annotations. Moreover, our method is generic. It can be applied to any CNN- or transformer-based deep classifier without the need for architectural change or adding significant training time. Our extensive evaluation [^1] Code: https://github.com/sbelharbi/interpretable-fer-aus. on two public benchmarks RAF-DB, and AffectNet datasets shows that our proposed strategy can improve layer-wise interpretability without degrading classification performance. In addition, we explore a common type of interpretable classifiers that rely on Class-Activation Mapping methods (CAMs), and we show that our training technique improves the CAM interpretability.
@@ 1 Introduction
Facial expression recognition (FER) has recently received much interest in the computer vision and machine learning communities [8, 82, 92]. Since facial expression is one of the most important ways for people to express emotions [16], FER finds a wide range of applications in, e.g., medical analysis and monitoring [1, 35, 84], eHealth [68], driver fatigue detection [2, 46], safe driving [33], security [41, 58], lecturing [69], and many other fields [63].
Figure 1: Comparison between standard non-interpretable and our interpretable FER classifier at inference time. Our classifier yields attention and CAM that are aligned with the expert’s knowledge used to assess basic facial expressions in images [51], which is illustrated in the map ${\bm{A}}$. Consequently, our model’s decision is more reliable. This is achieved without compromising the classification accuracy of our model. Details of our training strategy is presented in Fig.3. Note that in CAM-based models the classification head can be a fully convolutional module or standard fully connected layers that pooled posterior probabilities.
@@
FER remains challenging, particularly for in real world applications. This is due to the subtle differences between expressions, leading to low inter-class variability. As a result, learning to separate classes becomes difficult. This sets apart the FER task from standard classification problems, where classes are typically easy to distinguish. Additionally, people express the same facial expression differently depending on variant attributes such as level of expressiveness, age, gender, and ethnic background [38, 71]. To tackle these challenges, many deep methods have been proposed to build highly accurate models [8, 38, 82, 92]. This is achieved by learning deep, robust feature representations compared to traditional hand-crafted features [15, 47, 67]. Deep FER methods usually rely on either global or local approaches to learn feature representations. Global approaches [24, 39] propose training losses to improve representation robustness. To avoid incorporating noise into the global image representation, local approaches resort to learning from parts of the image [8, 30, 40, 74, 79, 80, 82, 92]. These methods use facial landmarks to extract robust features around relevant parts of the face [30, 79, 92]. Other methods rely on self-attention mechanisms to focus on relevant and discriminative parts of the image and suppress noisy regions [8, 40, 74, 80, 82].
@@
State-of-the-art FER methods for classification have achieved great progress in term of accuracy. However, in multiple critical applications such as eHealth [17, 61, 94], end-users may not require only an accurate classifier that yields a classification score. They may be also interested in an interpretable decision of the model [17, 83, 94]. For instance, this can help clinicians and therapists understand and build trust in the FER model’s decisions, planning better subsequent future interventions [72]. Subsequently, this facilitates better integrating machine learning-based models into daily clinical routines and health care practice [72]. In addition, interpretability can greatly help diagnose errors made by machine learning models and facilitate the identification of weaknesses for future improvement. Unfortunately, interpretability has been overlooked and shadowed in the FER task due to the main focus on classification accuracy. This has led to the design of FER models that lack interpretability. Recent attention-based works such as APVIT [82] build explicitly internal attention to select only discriminative regions of interest (ROIs). Although they provide visual interpretability to highlight the used regions for the classification, such ROIs are not necessarily aligned with what experts typically use to assess facial expressions. Commonly, experts rely on a codebook of action units to determine a basic facial expression [51] (Fig. 2). Each basic facial expression is associated with a subset of spatial action units in the face. Ideally, an interpretable classifier should point to and localize the correct action units’ ROIs when predicting a facial expression. However, this is a challenging task since predicting such ROIs using only image expression supervision is not trivial. Additionally, these localization cues are not publicly available to be learned since they burden the annotation cost and complicate building large-scale benchmarks.
@@
Figure 2: Codebook of basic facial expressions and their associated action units [51]. The spatial action units map is built using the image expression to select the right corresponding action unit subset in combination with facial landmarks which are employed to localize these same action units in the image properly. The code ’AUx’ is the identifier of the action unit [51].
@@
This work proposes a generic learning strategy to build an interpretable deep classifier for the FER task. This is achieved using spatial action units cues constructed from the image class supervision without extra manual annotation cost. This explicitly integrates the expert approach for facial expression assessment into the decision-making of a classifier. In particular, we build a visual interpretability tool consistent with action units cues used by experts [51]. To this end, we construct a spatial discriminative heatmap that gathers relevant locations of action units required to determine the expression in an input image. Then, layer-wise deep features of a classifier are constrained to be correlated with such heatmap cues during training. At the same time, the model is trained to correctly classify the image. Such a multi-task training scheme allows to build accurate but, most importantly, interpretable classifiers. During inference, the layer-wise attention plays the role of a visual interpretability tool that indicates which ROIs are used to determine the predicted expression (Fig. 3, 1). This is achieved without any extra manual annotation cost, as illustrated in Fig. 3. Discriminative action units heatmap can be built simply using image class expression, action units codebook, and facial landmarks, commonly estimated using off-the-shelf methods. Note that typically, interpretability in models usually emerges as a result of other tasks, such as classification without explicit supervision for interpretability [19]. It comes as a tool to clarify the model decision, for instance, under the form of visual ROIs. However, since we explicitly provide interpretability cues to train our model, we refer to it as guided interpretable FER model.
@@
Following the interpretable classifier direction, we explore a curated branch of classifiers in this work that allows us to perform image classification and yield visual interpretability. In particular, we visit Class-activation maps (CAMs)-based classifiers [12, 62]. Such models are popular in the weakly-supervised object localization task (WSOL) in computer vision for different imaging types, including natural scene [93] and medical imaging [62, 73]. Given an input image, such classifiers can be trained using only image class labels to correctly classify an image and provide a per-class heatmap, i.e., CAM, to localize ROIs related to the image class. Therefore, they play an important role in localization and interpretability, making them well-suited for our focus in this work. Therefore, we conduct a comparative study of different CAM-based methods from different families and assess their classification and interpretability capacity for the FER task. We note that this is the first work to leverage CAM-based methods for the FER task.
@@
(1) Unlike standard non-interpretable FER classifiers, we propose a generic learning strategy to build an accurate but, most importantly, interpretable deep classifier (Fig.1). The interpretability is consistent and aligned with the process used by experts to determine basic facial expressions, which relies on spatial action units.
(2) Our method does not: add any extra manual annotation, add significant extra computations during training, or change the model architecture or the inference process. In addition, our method is generic. It can be used with any deep CNN or transformer-based model.
(3) Our approach is validated over two public FER benchmarks: RAF-DB, and AffectNet by quantifying both classification and interpretability accuracy. We evaluated different CAM-based methods with and without our spatial action units cues. Different ablations are provided as well.
(4) Empirical results showed that both classification and interpretability improve with our method. Interestingly, compared to a vanilla deep classifier, we show that its layer-wise attention can largely be shifted using our action units’ spatial cues without compromising classification accuracy. In fact, we show that classification accuracy improves, particularly over large datasets such as AffectNet. This demonstrates that, indeed, spatial action units are a reliable source for discriminative ROIs for basic facial expression recognition [51].
@@ 2 Related Work
This review covers related work on FER classifier systems, CAM-based methods for localization and interpretability, interpretability in FER systems, and finally, action units in FER task.
Standard FER classification methods. With the availability of large public benchmarks [39, 53], different methods have been proposed aiming to achieve state-of-the-art classification accuracy [38]. This is most achieved by designing robust features to overcome the limitations of traditional hand-crafted features [15, 47, 67]. Building robust features is tackled by deep models either via a global or local approach.
@@
A second subset exploits attention mechanism [8, 80, 74, 81, 82] either self-learned or with provided cue. DAM-CNN [80] learns a spatial feature attention to filter out noise and keep only relevant spatial features to build a dense embedding later. The APVIT method [81, 82] leverages transformers and their attention potential to design an FER classifier. In particular, they propose a self-attention approach early in the network. Such attention allows the selection of relevant patch regions and performs hard attention at the spatial feature level. Patches are scored, and only the top-k are allowed to proceed into the next layer to build an image embedding. This explicitly allows the model to learn to filter out irrelevant patches. Other methods are provided external spatial cues to be aware where to look for discriminative regions [8, 40, 74]. RAN [74] is a region-based method. It performs image cropping either randomly or using landmarks. Their features are then attended using self-attention and relation-attention modules. Such attention is then used to combine different features and build an image embedding. This makes the final representation robust to pose and occlusion. Li et al. [40] follow a similar direction by decomposing the spatial features into the regions and using the attention module afterward. Bonnard et al. [8] use facial landmarks to build a heatmap, which is used to align spatial features in a deep model. This is expected to filter out irrelevant spatial features and focus mostly on features around landmarks. Our work is more aligned with this family of methods. However, we use grounded and more reliable cues, that is, spatial action units. Since spatial action units are class-dependent, they are more discriminative for the FER task [51], compared to facial landmarks, which are generic and class-agnostic.
@@
While the aforementioned methods focus on FER over still images, other methods tackle video applications [42, 43, 44, 45]. These methods deal with similar issues presented in still images while leveraging temporal dependency and multi-modality in videos to capture better expressiveness and build robust features.
@@
CAM-based methods for weakly-supervised object localization and interpretability. Using only image class as supervision, CAM-based methods can be trained to classify an image correctly and yield reliable localization of ROIs related to the image class. They are currently the dominant approach for the WSOL task [12, 62]. They achieved interesting results in different domains, including natural scene images [7, 62, 77], and medical imaging [5, 6, 62, 73]. They have been extended to localization in videos as well  [3, 4]. Early works have focused on building different spatial pooling layers [20, 21, 59, 93]. However, CAMs tend to highlight only small and most discriminative parts of an object [12, 62], limiting its localization performance. To overcome this, different strategies have been proposed, including data augmentation on input image or deep features [6, 13, 49], as well as architectural changes [29, 36, 77, 90]. Recently, learning via pseudo-labels has shown great potential, despite its reliance on noisy labels [7, 5, 52, 56, 54, 55, 75]. Most previous methods used only forward information in a CNN (bottom-up family [62]). However, other methods aim to leverage backward information as well (top-down methods [62]). These include biologically inspired methods [9, 87], and gradient [10, 28, 34, 66] or confidence score aggregation methods [18, 57]. CAM-based methods have been used for interpretability as well [64, 78] since they provide a map that indicates ROIs relevant to the model’s decision. Gradient-based CAMs are most common in interpretability task [10, 28, 34, 66] in addition to [9, 87]. They mainly search for ROIs in the feature maps that better stimulate a class response. This inspires other methods [88] to leverage gradient and extend it to input image and perform perturbation analysis [14, 25, 26, 60, 86] where the aim is to find which ROI of the network’s input that better stimulates its output. Compared to CAM-based, these methods are often expensive in term of computation and require performing an optimization, making them less attractive for FER tasks. However, CAM-based methods are simple to use, and they are built into the model, requiring a single forward or a forward and backward computation.
@@
Interpretability in FER systems. In critical FER applications such as eHealth and behavioral health interventions and assessments [11], it may not enough to simply accurately predict the facial expression. Users may require interpretability to help understand the model decision [17, 61, 83, 94]. Although important, interpretability in FER systems has been largely overlooked. This is mainly due to the sole focus on achieving state-of-the-art classification accuracy over challenging benchmarks [38]. The absence of public datasets with interpretability annotation has also contributed to shadowing such an important task. Some of the recent works have attempted to provide built-in discriminative attention under the form of visual interpretability, such as the APVIT method [81, 82]. These methods are still limited and, most importantly, unreliable since their interpretability has not been quantified yet due to the lack of annotation. In this work, we propose a protocol to evaluate the interpretability of an FER system without the need for extra manual annotation. We resort to using spatial facial action units to simulate experts’ processes to assess expressions [51]. Such action units are built and encoded automatically in a convenient spatial heatmap using only image expression as supervision. This can help to automatically label large benchmarks easily, allowing training and evaluation of FER models in term of interpretability.

### v3

@@ Abstract
Although state-of-the-art classifiers for facial expression recognition (FER) can achieve a high level of accuracy, they lack interpretability, an important feature for end-users. Experts typically associate spatial action units (AUs) from a codebook to facial regions for the visual interpretation of expressions. In this paper, the same expert steps are followed. A new learning strategy is proposed to explicitly incorporate AU cues into classifier training, allowing to train deep interpretable models. During training, this AU codebook is used, along with the input image expression label, and facial landmarks, to construct a AU heatmap that indicates the most discriminative image regions of interest w.r.t the facial expression. This valuable spatial cue is leveraged to train a deep interpretable classifier for FER. This is achieved by constraining the spatial layer features of a classifier to be correlated with AU heatmaps. Using a composite loss, the classifier is trained to correctly classify an image while yielding interpretable visual layer-wise attention correlated with AU maps, simulating the expert decision process. Our strategy only relies on image class expression for supervision, without additional manual annotations. Our new strategy is generic, and can be applied to any deep CNN- or transformer-based classifier without requiring any architectural change or significant additional training time. Our extensive evaluation [^1] Code :https://github.com/sbelharbi/interpretable-fer-aus. on two public benchmarks RAF-DB, and AffectNet datasets shows that our proposed strategy can improve layer-wise interpretability without degrading classification performance. In addition, we explore a common type of interpretable classifiers that rely on class activation mapping (CAM) methods, and show that our approach can also improve CAM interpretability.
@@
Figure 1: Comparison of class activation mapping (CAM) and attention maps produced at inference time using a FER classifier trained without (top) and with (bottom, ours) AU maps. In our experiments, we use an identical architecture to compare the impact of training with and without AU maps. The difference only resides in an additional AU-based training loss using AU maps. Training a FER classifier with our AU maps yields attention and CAM that are aligned with the expert’s knowledge used to assess basic facial expressions in images [51], as illustrated in the AU map ${\bm{A}}$. Consequently, our approach allows training a classifier that provides reliable interpretability, without compromising the classification accuracy. Details of our training strategy are presented in Fig.3. Note that in CAM-based models, the classification head can be fully convolutional or standard fully connected layers pooling posterior probabilities.
@@ 1 Introduction
Facial expression recognition (FER) has recently received much interest in the computer vision and machine learning communities [8, 82, 92]. Since facial expression is one of the most important ways for people to express emotions [16], FER finds a wide range of applications in, e.g., medical analysis and monitoring [1, 35, 84], e-health [68], driver fatigue detection [2, 46], safe driving [33], security [41, 58], lecturing [69], and many other fields [63].
@@
FER remains challenging, particularly for real-world applications. This is due to the subtle differences between expressions, leading to low inter-class variability. Learning to distinguish samples from overlapping classes can be difficult. Additionally, people express the same facial expression differently depending on attributes such as level of expressiveness, age, gender, and ethnic background [38, 71]. To tackle these challenges, many deep learning methods have been proposed to train highly accurate classification models [8, 38, 82, 92]. This is achieved by learning robust feature representations compared to traditional hand-crafted features [15, 47, 67]. Deep FER methods usually rely on either global or local approaches to learn feature representations. Global approaches [24, 39] propose training loss functions to improve the overall representation robustness. To avoid incorporating noise into the global image representation, local approaches resort to learning from parts of the image [8, 30, 40, 74, 79, 80, 82, 92]. These methods use facial landmarks to extract robust local features around relevant facial parts [30, 79, 92]. Other methods rely on self-attention mechanisms to focus on relevant and discriminative parts of the facial image and suppress noisy regions [8, 40, 74, 80, 82].
@@
State-of-the-art FER classifiers have achieved significant progress in terms of accuracy. However, end-users may not require only an accurate classifier that yields a classification score in multiple critical applications [17, 61, 94]. They may also need the model to provide an interpretable decision [17, 83, 94]. For instance, this can help clinicians and therapists understand and build trust in the FER model decisions, planning better future interventions [72]. Subsequently, this facilitates better integration of machine learning methods into daily clinical routines and health care practice [72]. In addition, interpretability can greatly help diagnose errors made by machine learning models and facilitate the identification of weaknesses for future improvement. Unfortunately, interpretability has been overlooked in the FER due to the main focus on classification accuracy. This has led to the design of FER models that lack interpretability. Recent progress with attention-based methods like APVIT [82] allows extracting internal attention explicitly to select only discriminative regions of interest (ROIs). Although they provide visual interpretability to highlight the regions used for classification, such ROIs are not necessarily aligned with expert knowledge typically used to assess facial expressions.
Experts commonly rely on a codebook of AUs to determine a basic facial expression [51] (Fig. 2). Each expression is associated with a subset of spatial AUs in the face. Ideally, an interpretable classifier should point to and localize the correct AU ROIs when predicting the corresponding expression. However, this task is challenging since predicting such ROIs using only image expression supervision is not trivial. Additionally, the localization cues are not publicly available to be learned due to the annotation costs and the complexity of building large-scale benchmarks.
@@
Figure 2: Codebook of basic facial expressions and their associated AUs [51]. The spatial AU map is built using the image expression to select the right corresponding AU subset in combination with facial landmarks, which are employed to localize these same AUs in the image. In particular, the location of landmarks is used to estimate AU positions. For instance, the right ’Cheek’ location is estimated using landmark 47 (middle of the low right eye) and 11 (right side of the jaw). The code ’AUx’ is the identifier of the AU [51].
@@
This paper introduces a generic learning strategy to build an interpretable deep classifier for the FER task. This is achieved using spatial AU cues constructed from the image class supervision without extra manual annotation cost. This explicitly integrates the expert facial expression assessment approach into a classifier’s decision-making. In particular, we build a visual interpretability tool consistent with AU cues used by experts [51]. To this end, a spatial discriminative heatmap is constructed that gathers relevant locations of AUs required to determine the expression in an input image. Then, layer-wise deep features of a classifier are constrained to be correlated with such heatmap cues during training. At the same time, the model is trained to classify the image correctly. Such a multi-task training scheme allows for building accurate but, most importantly, interpretable classifiers. During inference, the layer-wise attention provides a visual interpretability tool that indicates which ROIs are used to determine the predicted expression (Fig. 3, 1). This is achieved without additional annotation costs, as illustrated in Fig. 3. A discriminative AU heatmap can be built simply using image class expression, an AU codebook, and facial landmarks, which are commonly estimated using off-the-shelf methods. Interpretability in models usually emerges as a result of other tasks, such as classification without explicit supervision for interpretability [19]. It represents a tool to clarify the model decision, for instance, under the form of visual ROIs. However, since we explicitly provide interpretability cues to train our model, we refer to it as guided interpretable FER model.
@@
Following the interpretable classifier direction, we explore a curated branch of classifiers in this work that allows us to perform image classification and yield visual interpretability. In particular, we employ CAM-based classifiers [12, 62]. These methods are popular for weakly-supervised object localization (WSOL) tasks in computer vision on different imaging types, including natural scene [93] and medical imaging [62, 73]. Given an input image, a classifier can be trained using only image-class labels to correctly classify an image and provide a per-class heatmap, i.e., CAM, to localize image ROIs related to the output prediction. Therefore, they play an important role in localization and visual interpretability, making them well-suited for our work. Consequently, we conduct a comparative study using different CAM methods from other families of methods, and assess their capacity for classification and interpretability on the FER task. Note that this is the first work to leverage CAM methods for the FER task.
@@
(1) To improve the visual interpretability of state-of-the-art FER classifiers, we introduce a learning strategy allowing to training of an accurate but, most importantly, interpretable deep classifier (Fig.1) that is consistent with the process used by experts to determine basic facial expressions. Our training relies on aligning spatial features with spatial AU maps built using facial landmarks and image-class labels. This guides the classifier’s attention to use regions around AUs leading to more interpretable decisions.
(2) Our method does not require extra manual annotation, significant extra computations during training, nor change the model architecture or the inference process. In addition, our method is generic – it can be used with any deep CNN or transformer-based model.
(3) Our approach is validated experimentally on two public FER benchmarks (RAF-DB, and AffectNet) in terms of classification and interpretability accuracy. We evaluated different CAM-based methods with and without our spatial AU cues. Different ablations are provided as well. Empirical results showed that both classification and interpretability improved. Interestingly, compared to a vanilla deep classifier, we show that its layer-wise attention can largely be shifted using our AU spatial cues without compromising classification accuracy. We show that classification accuracy improves, particularly over large datasets such as AffectNet. This demonstrates that spatial AUs are a reliable source for discriminative ROIs for basic facial expression recognition [51].
@@ 2 Related Work
This review covers related work on FER classifier systems, CAM methods for localization and interpretability, FER systems interpretability, and AUs in FER tasks.
2.1 Standard FER classifiers
Given the availability of large public benchmarks [39, 53], different methods have been proposed to achieve state-of-the-art classification accuracy [38]. This is most achieved by designing robust features to overcome the limitations of traditional hand-crafted features [15, 47, 67]. Robust features are also learned using deep models, either through a global or local approach.
@@ 2.1 Standard FER classifiers
A second subset exploits attention mechanism [8, 80, 74, 81, 82] either self-learned or with provided cue. DAM-CNN [80] learns a spatial feature attention to filter out noise and keep only relevant spatial features to build a dense embedding later. The APVIT method [81, 82] leverages transformers and their attention potential to design a FER classifier. In particular, they propose a self-attention approach early in the network. Such attention allows the selection of relevant patch regions and performs hard attention at the spatial feature level. Patches are scored, and only the top-k are allowed to proceed into the next layer to build an image embedding. This explicitly allows the model to learn to filter out irrelevant patches. Other methods are provided external spatial cues to be aware where to look for discriminative regions [8, 40, 74]. RAN [74] is a region-based method. It performs image cropping either randomly or using landmarks. Their features are then attended using self-attention and relation-attention modules. Such attention is then used to combine different features and build an image embedding. This makes the final representation robust to pose and occlusion. Li et al. [40] follow a similar direction by decomposing the spatial features into the regions and using the attention module afterward. Bonnard et al. [8] use facial landmarks to build a heatmap, which is used to align spatial features in a deep model. This is expected to filter out irrelevant spatial features and focus mostly on features around landmarks. Our work is more aligned with this family of methods. However, we use grounded and more reliable cues, that is, spatial AUs. Since spatial AUs are class-dependent, they are more discriminative for the FER task [51], compared to facial landmarks, which are generic and class-agnostic.
@@
While the abovementioned methods focus on FER over still images, other methods tackle video applications [42, 43, 44, 45]. These methods deal with similar issues presented in still images while leveraging temporal dependency and multi-modality in videos to capture better expressiveness and build robust features.
2.2 CAM methods for WSOL and interpretability
@@
Using only image class as supervision, CAM-based methods can be trained to classify an image correctly and yield reliable localization of ROIs related to the image class. They are currently the dominant approach for the WSOL task [12, 62]. They achieved interesting results in different domains, including natural scene images [7, 62, 77], and medical imaging [5, 6, 62, 73]. They have also been extended for localization in videos [3, 4]. Early works have focused on building different spatial pooling layers [20, 21, 59, 93]. However, CAMs tend to highlight only small and most discriminative parts of an object [12, 62], limiting its localization performance. To overcome this, different strategies have been proposed, including data augmentation on input image or deep features [6, 13, 49], as well as architectural changes [29, 36, 77, 90].
Recently, learning via pseudo-labels has shown great potential, despite its reliance on noisy labels [7, 5, 52, 56, 54, 55, 75]. Most previous methods used only forward information in a CNN (bottom-up family [62]). However, other methods also aim to leverage backward information (top-down methods [62]). These include biologically inspired methods [9, 87], and gradient [10, 28, 34, 66] or confidence score aggregation methods [18, 57]. CAM-based methods have been used for interpretability as well [64, 78] since they provide a map that indicates ROIs relevant to the model’s decision. Gradient-based CAMs are most common in interpretability task [10, 28, 34, 66] in addition to [9, 87]. They mainly search for ROIs in the feature maps that better stimulate a class response. This inspires other methods [88] to leverage gradient and extend it to input image and perform perturbation analysis [14, 25, 26, 60, 86] where the aim is to find which ROI of the network’s input that better stimulates its output. Compared to CAM-based, these methods are often more expensive in computation and require optimization, making them less attractive for FER tasks. However, CAM-based methods are simple to use, and they are built into the model, requiring a single forward or a forward and backward computation.
2.3 Interpretability in FER systems
@@
In critical FER applications such as e-health and behavioral health interventions and assessments [11], it may not be enough to predict the facial expression accurately. Users may require interpretability to help understand the model decision [17, 61, 83, 94]. Although important, interpretability in FER systems has been largely overlooked. This is mainly due to the focus on achieving state-of-the-art classification accuracy over challenging benchmarks [38]. The absence of public datasets with interpretability annotation has also contributed to shadowing such an important task. Some recent works have attempted to provide built-in discriminative attention under visual interpretability, such as the APVIT method [81, 82]. Due to the lack of annotation, these methods are still limited and unreliable since their interpretability has not yet been quantified. In this work, we propose a protocol to evaluate the interpretability of a FER system without the need for extra manual annotation. We use spatial facial AUs to simulate experts’ processes to assess expressions [51]. Such AUs are built and encoded automatically in a convenient spatial heatmap using only image expression as supervision. This can help to automatically label large benchmarks easily, allowing training and evaluation of FER models in terms of interpretability.
2.4 AUs in FER systems